# Chaos Engineering: A Multi-Vocal Literature Review

### v1

@@ Abstract.
Organizations, particularly medium and large enterprises, typically today rely heavily on complex, distributed systems to deliver critical services and products. However, the growing complexity of these systems poses challenges in ensuring service availability, performance, and reliability. Traditional resilience testing methods often fail to capture modern systems’ intricate interactions and failure modes. Chaos Engineering addresses these challenges by proactively testing how systems in production behave under turbulent conditions, allowing developers to uncover and resolve potential issues before they escalate into outages. Though chaos engineering has received growing attention from researchers and practitioners alike, we observed a lack of a comprehensive literature review. Hence, we performed a Multivocal Literature Review (MLR) on chaos engineering to fill this research gap by systematically analyzing 88 academic and grey literature sources published from January 2019 to April 2024. We first used the selected sources to derive a unified definition of chaos engineering and to identify key capabilities, components, and adoption drivers. We also developed a taxonomy for chaos engineering and compared the relevant tools using it. Finally, we analyzed the state of the current chaos engineering research and identified several open research issues.
@@
copyright: noneisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
An effective approach to overcoming the challenges of ensuring reliable and resilient cloud-based systems is chaos engineering, which proactively tests system resilience by introducing controlled failures into production environments (Bhaskar, 2022; Fawcett, 2020; Palani and Gupta, 2023). This approach was pioneered by Netflix in 2008 to fortify its video streaming infrastructure. Since then, chaos engineering has grown and extended its influence, evolving from its original focus on cloud services and microservice architecture to broader applications across industries (Bairyev, 2023). Today, this approach is widely employed in critical sectors like fintech and healthcare, with major technology companies such as LinkedIn, Facebook, Google, Microsoft, Amazon, Slack, and others adopting it to enhance their system resilience, improve customer experience, and drive business growth.  (Inc., 2023; Patel, 2022; Roa, 2022; Fawcett, 2020; Al-Said Ahmad et al., 2024; Bremmers, 2021; Bairyev, 2023).
Figure 1 illustrates chaos engineering’s adoption through Moore’s ’Crossing the Chasm’ model (Moore and McKenna, 1999). Adoption begins with Innovators (early risk-takers) and Early Adopters (those seeking competitive advantage), followed by the Early Majority (who adopt once proven), the Late Majority (who wait until widely supported), and finally, Laggards (the last to adopt, often only when essential). The critical hurdle, Moore’s Chasm, is where many approaches fail due to complexity and unclear benefits. Crossing this chasm leads to mainstream adoption and integration across the SDLC (Software Development Lifecycle) and industries (Mukkara, 2022).
Figure 1. Chaos Engineering in the Context of Geoffrey Moore’s Technology Adoption Lifecycle Model from Crossing the Chasm  (Mukkara, 2022).
[…]
Chaos Engineering on the Technology Adoption Lifecycle
@@
We observed many researchers and practitioners have explored chaos engineering in various contexts, discussing its implementations and challenges through research papers and grey literature such as online blogs, technical papers, and forums. However, despite its increasing adoption, much of the existing knowledge remains dispersed across the literature. To the authors’ best understanding, no secondary research has been published to synthesize this scattered information, such as comprehensive reviews or survey papers. Hence, to address this research gap, in this paper, we present a Multivocal Literature Review (MLR) (Garousi et al., 2019) on chaos engineering, synthesizing a broad spectrum of perspectives from academic and grey literature sources. Unlike traditional Systematic Literature Reviews (SLRs), which often overlook grey literature, an MLR includes insights from peer-reviewed academic studies alongside industry reports, technical blogs, and white papers (Garousi et al., 2019, 2016). In particular, as chaos engineering emerged from the industry and most tools are from the industry, the grey literature is a highly relevant and valuable source of information.
In this MLR, we selected and analyzed 88 sources, spanning academic and grey literature sources. Our findings provide a detailed overview of chaos engineering, including its core principles, quality requirements, and foundational components defined by researchers and industry experts. We also examine the motivations behind adopting chaos engineering, the specific resilience challenges it addresses, and the practices and evaluation methods that support its effective implementation. Additionally, we introduce a taxonomy that can help practitioners and decision-makers make informed decisions when comparing and selecting chaos engineering tools. By analyzing the benefits and shortcomings of current approaches in the chaos engineering domain, this MLR contributes to a thorough understanding of chaos engineering and provides guidelines for future studies to advance this critical field.
@@
The remainder of the paper is organized as follows. Section  2 presents the research methodology used in this study, while Section 3 provides a comprehensive definition of chaos engineering, along with its functionalities, quality requirements, and core components. Section 4 focuses on the motivations for chaos engineering and the specific challenges it addresses. Section 5 introduces a taxonomy of chaos engineering, covering tools, practices, and evaluation methods within the field. Section 6 examines the current state of chaos engineering in research and industry. Section  7 discusses the implications of our findings, highlighting key insights, identifying open issues, and addressing potential threats to validity. Finally, Section 8 summarizes the related literature surveys, and Section  9 concludes the paper.
@@ 2. Research Methodology
We employed the guidelines proposed by Garousi et al. (Garousi et al., 2019) to systematically collect and analyze multivocal literature on chaos engineering. We also drew methodological inspiration from other multivocal literature reviews, such as Buck et al. (Buck et al., 2021) and Islam et al. (Islam et al., 2019). Figure  2 illustrates our MLR methodology, comprising several steps divided into three phases: planning and designing the review protocol, executing the review, and documenting the findings. The remainder of this section provides a detailed discussion of each step.
@@ 2.1. Research Identification
We employed a search strategy to identify appropriate literature grounded in a series of research questions in Table  1. The first three questions cover the various aspects of chaos engineering, including its definitions, key features, benefits, challenges, use cases, tools and techniques, and best and bad practices. The final question provides an overview of the present landscape of chaos engineering literature.
@@
RQ1. How is chaos engineering defined in the literature? RQ1.1. Which core principles and primary functionalities are most commonly associated with chaos engineering? RQ1.2. What are the core components of a chaos engineering platform, and how do they support its core functions? RQ1.3. What quality requirements are essential for an effective chaos engineering platform, and how are they quantified? This question clarifies the definition of chaos engineering, exploring its core principles, key components, and quality requirements. These elements build a foundation for understanding, implementing, and evaluating chaos engineering solutions.
@@ 2.2.1. Sources of data
We conducted automated searches in four major digital libraries—Scopus, IEEE Xplore, ACM Digital Library, and Web of Science—using both peer-reviewed scholarly articles and grey literature as sources. These libraries serve as primary sources for software engineering studies (Chen et al., 2010). We performed multi-keyword searches targeting titles, abstracts, and keywords for papers published between January 2019 and April 2024. Additionally, Google Scholar was used to pinpoint any relevant literature that was not included in the previous searches. We used Google Search for grey literature, similar to other multi-vocal literature reviews (MLRs)  (Garousi and Mäntylä, 2016; Garousi et al., 2017; Islam et al., 2019).
@@ 2.3. Eligibility Criteria
Following some existing MLR studies (Islam et al., 2019; Butijn et al., 2020; Garousi et al., 2019), we applied a specific set of inclusion and exclusion criteria, as detailed in Table  2
@@
IC2: Articles must include robust validation methods, including working prototypes or tools with proper references, especially for grey sources. EC2: Articles that lack the relevant focus on chaos engineering.
@@ 2.4. Study Selection
The libraries’ advanced query modes were used to maximize search coverage. The search was conducted in April 2024, and the results were saved in BibTeX and CSV formats. Duplicates were removed before proceeding with further analysis. Figure  3 illustrates the selection process for academic and grey literature, showing the databases searched and the number of selected papers at each stage. Different approaches were followed to select grey and academic literature.
@@ 2.4.1. Academic Literature Selection
Searches in Scopus, IEEE Xplore, ACM DL, and Web of Science yielded 1,290 papers. After duplicate removal, 222 unique papers remained. Applying the inclusion and exclusion criteria, the selection was narrowed to 42 relevant papers. An additional search was conducted using Google Scholar, which added two more papers to the final count, bringing the total to 44 papers.
Figure 3. MLR Study Selection Process.
@@
Criteria Questions
@@
Date (Q6). Is the publication date of the source within the past five years?
Table 4. Search Strings and Results.
Search Strings ACM DL IEEE Xplore Scopus Web of Science Total
"chaos engineering" OR "chaos test" OR "chaos experiment" OR "chaos toolkit" OR "chaos platform" OR "chaos mesh" OR "chaos integration" OR "chaos architecture" OR "chaos infrastructure" OR "chaos injection" OR "chaos tool" 95 63 121 47 326
"chaos engineering" AND ("trends" OR "developments" OR "progression") 29 0 17 0 46
"chaos engineering" AND ("leader" OR "pioneer" OR "expert" OR "contributor") 35 3 4 0 42
"chaos engineering" AND ("phases" OR "stages" OR "steps" OR"pipeline") 62 5 11 0 78
"chaos engineering" AND ("challenges" OR "issues" OR "problems") 79 16 34 4 133
"chaos engineering" AND ("applications" OR "benefits") 77 20 37 8 142
"chaos engineering" AND ("solutions" OR "approaches" OR "methods") 72 20 55 4 151
"chaos engineering" AND ("tools" OR "techniques") 79 24 36 8 147
"chaos engineering" AND ("best practices" OR "bad practices" OR "implementation" OR "Culture") 68 8 16 5 97
"chaos engineering" AND ("evaluation" OR "metrics") 75 24 26 3 128
Total 671 183 357 79 1290
@@ 2.4.2. Grey Literature Selection
We followed the approach of other multi-vocal and grey literature reviews  (Garousi and Mäntylä, 2016; Garousi et al., 2017; Islam et al., 2019; Butijn et al., 2020; Garousi et al., 2019) and used Google Search to gather grey literature. As described by these reviews, we examined the first ten pages of results, as this was deemed sufficient to identify the most pertinent literature since Google’s algorithm emphasizes the most relevant results on the first pages  (Soldani et al., 2018; Verdecchia et al., 2019; Islam et al., 2019). We continued scanning further only if necessary, stopping when no relevant new articles appeared  (Islam et al., 2019). To avoid bias, we searched in incognito mode, following the recommendations of Rahman et al. (Rahman et al., 2023). We incorporated additional quality assessment criteria beyond the inclusion and exclusion criteria, as outlined in Table  3. We picked six key criteria: the credibility of the publishing organization (Q1), the author’s prior publications in the field (Q2), the author’s institutional affiliation (Q3), clarity of purpose (Q4), the author’s expertise (Q5), and recency of the source (Q6). The responses to these criteria were classified as ‘Yes’ (1), ‘Partly’ (0.5), and ‘No’ (0), with scores assigned accordingly for each category. These criteria were adapted from prior studies on grey and multi-vocal literature reviews  (Soldani et al., 2018; Butijn et al., 2020; Kumara et al., 2021), ensuring a thorough evaluation of the sources. As a result, we selected 44 documents, contributing to 88 unique articles (Figure  3) for extracting and synthesizing data. Table  5 illustrates the selected sources’ distribution across various venues.
@@
Websites & Blogs Whitepaper
(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Zhang et al., 2021a; Cotroneo et al., 2022; Camacho et al., 2022; Torkura et al., 2021; Poltronieri et al., 2022; Zhang et al., 2023; Nikolaidis et al., 2021; Frank et al., 2021b, a; Pierce et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Chen et al., 2024; Nikolaidis et al., 2023; Ma et al., 2023; Meiklejohn et al., 2021; Frank et al., 2023; Sondhi et al., 2021; Kassab et al., 2022; Jernberg et al., 2020; Konstantinou et al., 2021; Pulcinelli et al., 2023; Ji et al., 2023; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Park et al., 2019; Klinaku et al., 2022; Rivera et al., 2023; Camilli et al., 2022; Yu et al., 2021; Siwach et al., 2022) = 44 (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Hawkins, 2020; Nombela, 2023; Le, 2022; National Australia Bank, 2020; Shah, 2021; Krishnamurthy, 2021; Moon, 2022; Katirtzis, 2022; Wachtel, 2023; Durai, 2022; Anwar and Arunachalam, 2019; Tomka, 2024; Dua, 2024; Bocetta, 2019; Mooney, 2023; Wu, 2021; Lardo, 2019; Miles, 2019a; Bhaskar, 2022; Azure Readiness, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Sbiai et al., 2023; Abdul, 2024; Roa, 2022; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Vinisky, 2024; Colyer, 2019) = 41 (Palani and Gupta, 2023; Tivey et al., 2019; Starr et al., 2022) = 3

### v2

@@ Abstract.
Organizations, particularly medium and large enterprises, typically rely heavily on complex, distributed systems to deliver critical services and products. However, the growing complexity of these systems poses challenges in ensuring service availability, performance, and reliability. Traditional resilience testing methods often fail to capture the intricate interactions and failure modes of modern systems. Chaos Engineering addresses these challenges by proactively testing how systems in production behave under turbulent conditions, allowing developers to uncover and resolve potential issues before they escalate into outages. Though chaos engineering has received growing attention from researchers and practitioners alike, we observed a lack of reviews that synthesize insights from both academic and grey literature. Hence, we conducted a Multivocal Literature Review (MLR) on chaos engineering to address this research gap by systematically analyzing 96 academic and grey literature sources published between January 2016 and April 2024. We first used the chosen sources to derive a unified definition of chaos engineering and to identify key functionalities, components, and adoption drivers. We also developed a taxonomy for chaos engineering platforms and compared the relevant tools using it. Finally, we analyzed the current state of chaos engineering research and identified several open research issues.
@@
copyright: acmlicensedisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
An effective approach to overcoming the challenges of ensuring reliable and resilient cloud-based systems is chaos engineering, which proactively tests system resilience by introducing controlled failures into production environments (Bhaskar, 2022; Fawcett, 2020; Palani and Gupta, 2023a). This approach was pioneered by Netflix in 2008 to fortify its video streaming infrastructure. Since then, chaos engineering has grown and extended its influence, evolving from its original focus on cloud services and microservice architecture to broader applications across industries (Bairyev, 2023). Today, this approach is widely employed in critical sectors, such as fintech and healthcare, with major technology companies like LinkedIn, Facebook, Google, Microsoft, Amazon, and Slack adopting it to enhance system resilience, improve customer experience, and drive business growth.  (Inc., 2023; Patel, 2022; Roa, 2022; Fawcett, 2020; Al-Said Ahmad et al., 2024; Bremmers, 2021; Bairyev, 2023).
@@
We observed that many researchers and practitioners have explored chaos engineering in various contexts, discussing its implementations and challenges through research papers and grey literature such as online blogs, technical papers, and forums. However, despite its increasing adoption, much of the existing knowledge remains dispersed across the literature. To the authors’ best understanding, no secondary research has been published to synthesize this scattered information, such as comprehensive reviews or survey papers. Hence, to address this research gap, in this paper, we present a Multivocal Literature Review (MLR) (Garousi et al., 2019) on chaos engineering, synthesizing a broad spectrum of perspectives from academic and grey literature sources. Unlike traditional Systematic Literature Reviews (SLRs), which often overlook grey literature, an MLR includes insights from peer-reviewed academic studies alongside industry reports, technical blogs, and white papers (Garousi et al., 2019, 2016). In particular, as chaos engineering emerged from the industry and most tools are from the industry, the grey literature is a highly relevant and valuable source of information. In this MLR, we selected and analyzed 96 sources, encompassing both academic and grey literature. We selected 2016 as the starting point, when chaos engineering principles were first formally articulated by Basiri et al (Basiri et al., 2016). Our findings provide a detailed overview of chaos engineering, including its core principles, quality requirements, and foundational components defined by researchers and industry experts. We also examine the motivations behind adopting chaos engineering, the specific resilience challenges it addresses, and the practices and evaluation methods that support its effective implementation. Additionally, we introduce a taxonomy of chaos engineering platforms and their capabilities, designed to support informed decision-making around tool adoption and deployment strategies. By analyzing the benefits and shortcomings of current approaches in the chaos engineering domain, this MLR contributes to a thorough understanding of chaos engineering and provides guidelines for future studies to advance this critical field.
@@
The remainder of the paper is organized as follows. Section 2 outlines the background and motivation for this review, including a discussion of related surveys, the distinction between chaos engineering and traditional testing, and the need to extend the existing body of knowledge on chaos engineering. Section 3 presents the research methodology used in this study, while Section 4 provides a comprehensive definition of chaos engineering, along with its functionalities, quality requirements, and core components. Section 5 focuses on the motivations for chaos engineering and the specific challenges it addresses. Section 6 introduces a taxonomy of chaos engineering, covering tools, practices, and evaluation methods within the field. Section 7 examines the current state of chaos engineering in research and industry. Section  8 discusses the implications of our findings, highlighting key insights, identifying open issues, and addressing potential threats to validity. Finally, Section  9 concludes the paper.
2. Background and Motivation
This section outlines the motivations for conducting this review by examining the limitations of existing literature (Section2.1), explaining how chaos engineering diverges from traditional testing methods (Section 2.2), and establishing the novelty and timeliness of this multivocal literature review (Section 2.3).
2.1. Related Work and Survey Gap
In this section, we consider key surveys on resilience, robustness, and fault injection techniques, exploring their relationship to the proactive focus of chaos engineering. For instance, Natella et al. (Natella et al., 2016) surveyed software fault injection techniques, categorizing faults by origin such as software, hardware, and environment, and by injection approach such as code changes or fault effect injection including state and interface errors. However, their focus is on pre-deployment environments (e.g., simulators, testbeds), whereas chaos engineering involves live, iterative experiments in production-like systems. Similarly, Colman-Meixner et al. (Colman-Meixner et al., 2016) focus on cloud resilience through redundancy and failover, but explore reactive recovery rather than proactive experimentation, as in chaos engineering. Laranjeiro et al. (Laranjeiro et al., 2021) consider static robustness techniques such as fuzzing and API testing, which lack real-time, adaptive feedback loops. Ruholamini et al.  (Rouholamini et al., 2020) surveyed self-healing in cloud-native systems, emphasizing failure detection and recovery rather than proactive failure exploration.
More recently, Arsecularatne and Wickramarachchi (Arsecularatne and Wickramarachchi, 2023) conducted a systematic review and qualitative study on the adoptability of chaos engineering in DevOps. Their work highlights organizational challenges and potential benefits but does not provide a structured analysis of tools, architectures, or evaluation metrics. Similarly, Mhatre et al. (Mhatre et al., 2024) examined chaos engineering within DevSecOps and SRE (Site Reliability Engineering) contexts, emphasizing conceptual alignment and use cases, but without offering a comprehensive taxonomy or resilience assessment frameworks. In summary, existing reviews examine fault injection or discuss chaos engineering within DevOps contexts. However, they offer little or no clear insight into its unified definition, functionalities, architectural components, quality attributes, comprehensive tool taxonomy, or the technical and organizational challenges of implementing it in production-like environments. This review addresses these gaps by presenting the first multivocal literature review (MLR) on chaos engineering, integrating both academic and industry sources to inform future research and practical adoption.
2.2. Chaos Engineering vs. Traditional Software Testing Techniques
Traditional software testing techniques, including unit, integration, stress, and robustness testing, primarily focus on validating correctness and performance under expected conditions in pre-deployment environments (Basiri et al., 2019; Shekhar, 2024). These approaches typically operate in isolated, controlled settings and verify that system components behave correctly with known inputs (Ji et al., 2023; Shekhar, 2024). However, in modern distributed, cloud-native systems, such methods often fail to expose cascading or emergent failures caused by asynchronous interactions and dynamic runtime behaviors (Migirditch et al., 2022; Ji et al., 2023; Fogli et al., 2023). Chaos engineering addresses these limitations by introducing controlled faults into production or production-like systems to explore system resilience (Zhang et al., 2019; Torkura et al., 2020; Fogli et al., 2023). Rather than validating functionality using predefined test scripts, chaos experiments simulate failures, such as service crashes, latency spikes, or regional outages,to evaluate system behavior and recovery under adverse runtime conditions (Zhang et al., 2019; Migirditch et al., 2022; Shekhar, 2024). This methodology helps surface hidden vulnerabilities that traditional test environments often miss (Ji et al., 2023; Parsons, 2021). Importantly, chaos engineering is not a replacement for conventional testing, but a complement. While traditional methods confirm system correctness under expected conditions, chaos engineering validates system behavior under known faults and real-world failure scenarios  (Basiri et al., 2019; Fogli et al., 2023). This paradigm shift—from pre-deployment pass/fail validation to post-deployment resilience assessment—is what makes chaos engineering essential for today’s complex, distributed infrastructure (Zhang et al., 2019; Ji et al., 2023; Fogli et al., 2023). To clarify these distinctions, Table 1 highlights key differences between traditional testing and chaos engineering. The comparison focuses on three main areas: the testing environment (pre-production vs. production-like systems), the primary objective (functional verification vs. uncovering hidden system weaknesses), and the types of failures addressed (expected vs complex real-world faults).
Table 1. Comparison Between Traditional Software Testing Techniques and Chaos Engineering
Category Traditional Testing Chaos Engineering Articles Referenced
Testing Environment Pre-production or simulated systems. Production or production-like systems. (Bairyev, 2023; Zhang et al., 2019; Simonsson et al., 2021; Zhang et al., 2021b; Gill, 2021; Bocetta, 2019; Bhaskar, 2022; Krivas and Portela, 2020; Wilms, 2018)
Testing Objective Verifies specific functions or components against pre-defined requirements and standards. Identifies obscure system weaknesses that may only become apparent during turbulent failure conditions. (Torkura et al., 2020; Inc., 2023; Bairyev, 2023)
Types of Failures Focuses on expected and known usage scenarios and errors to verify how the system behaves under normal or predictable conditions. Simulates real-world failures (e.g., network outages, high load, hardware issues) to see how the system reacts to these changes. (Torkura et al., 2020; Pierce et al., 2021; Ji et al., 2023; Konstantinou et al., 2021; Dua, 2024; Miles, 2019a; Bairyev, 2023)
2.3. Research Contribution and Novelty
Despite the growing popularity of chaos engineering (Monroe, 2019), there remains a lack of unified academic frameworks that synthesize its principles, classify its tool landscape, or integrate its practices across both research and industry domains. To the best of our knowledge, this paper presents the first Multivocal Literature Review (MLR) on chaos engineering, integrating both academic studies and grey literature. Our contributions are fourfold. First, we synthesize and unify definitions, core principles, functionalities, and architectural components of chaos engineering from academic and industry perspectives (RQ1). Second, we identify the key technical and socio-technical challenges driving its adoption (RQ2). Third, we propose a taxonomy of chaos engineering tools, classify evaluation methods, and highlight adoption practices (RQ3). Fourth, we analyze publication trends, venue types, and key contributors to assess the current state of research and practice (RQ4). Together, these contributions provide a foundational reference on chaos engineering that bridges research and practice, supports rigorous evaluation of chaos engineering tools, enables their systematic adoption in industry, and offers directions for future research.
@@ 3. Research Methodology
We employed the guidelines proposed by Garousi et al. (Garousi et al., 2019) to systematically collect and analyze multivocal literature on chaos engineering. We also drew methodological inspiration from other multivocal literature reviews, such as those by Buck et al. (Buck et al., 2021) and Islam et al. (Islam et al., 2019). Figure 1 illustrates our MLR methodology, comprising several steps divided into three phases: planning and designing the review protocol, executing the review, and documenting the findings. The remainder of this section provides a detailed discussion of each step.
@@ 3.1. Research Identification
We employed a search strategy to identify appropriate literature grounded in a series of research questions in Table 2. The first three questions cover the various aspects of chaos engineering, including its definitions, key features, benefits, challenges, use cases, tools and techniques, and best and bad practices. The final question provides an overview of the present landscape of chaos engineering literature.
@@
RQ1. How is chaos engineering defined in the literature? RQ1.1. Which core activities and primary functionalities are most commonly associated with chaos engineering? RQ1.2. What are the core components of a chaos engineering platform, and how do they support its core functions? RQ1.3. What quality requirements are essential for an effective chaos engineering platform, and how are they quantified? This question clarifies the definition of chaos engineering, exploring its core activities, key components, and quality requirements. These elements build a foundation for understanding, implementing, and evaluating chaos engineering solutions.
@@ 3.2.1. Sources of data
We conducted automated searches in four major digital libraries—Scopus, IEEE Xplore, ACM Digital Library, and Web of Science—using both peer-reviewed scholarly articles and grey literature as sources. These libraries serve as primary sources for software engineering studies (Chen et al., 2010). Following the formal articulation of chaos engineering principles by Basiri et al. in 2016 (Basiri et al., 2016), we limited our search to publications from January 2016 to April 2024. We performed multi-keyword searches over titles, abstracts, and keywords within this range. Additionally, Google Scholar was used to pinpoint any relevant literature that was not included in the previous searches. We used Google Search for grey literature, similar to other multi-vocal literature reviews (MLRs) (Garousi and Mäntylä, 2016; Garousi et al., 2017; Islam et al., 2019).
@@ 3.2.2. Search Terms
Table 3. Search Strings and Results.
Search Strings ACM DL IEEE Xplore Scopus Web of Science Total
"chaos engineering" OR "chaos test" OR "chaos experiment" OR "chaos toolkit" OR "chaos platform" OR "chaos mesh" OR "chaos integration" OR "chaos architecture" OR "chaos infrastructure" OR "chaos injection" OR "chaos tool" 108 81 142 57 388
"chaos engineering" AND ("trends" OR "developments" OR "progression") 30 0 20 0 50
"chaos engineering" AND ("leader" OR "pioneer" OR "expert" OR "contributor") 39 3 4 0 46
"chaos engineering" AND ("phases" OR "stages" OR "steps" OR "pipeline") 71 6 11 1 89
"chaos engineering" AND ("challenges" OR "issues" OR "problems") 91 19 37 4 151
"chaos engineering" AND ("applications" OR "benefits") 89 22 43 9 163
"chaos engineering" AND ("solutions" OR "approaches" OR "methods") 84 22 60 4 170
"chaos engineering" AND ("tools" OR "techniques") 92 30 42 10 174
"chaos engineering" AND ("best practices" OR "bad practices" OR "implementation" OR "Culture") 77 9 17 5 108
"chaos engineering" AND ("evaluation" OR "metrics") 85 25 28 4 142
Total 766 217 404 94 1481
@@ 3.3. Eligibility Criteria
Following some existing MLR studies (Islam et al., 2019; Butijn et al., 2020; Garousi et al., 2019), we applied a specific set of inclusion and exclusion criteria, as detailed in Table 4
@@
IC2: Articles must include clear validation or evaluation methods, with supporting references. EC2: Articles that lack the relevant focus on chaos engineering.
@@ 3.4. Study Selection
The libraries’ advanced query modes were used to maximize search coverage. The search was conducted in April 2024, and the search results were exported from each database in BibTeX and CSV formats for reference management and data processing. Duplicates were removed before proceeding with further analysis. Figure  2 illustrates the selection process for academic and grey literature, showing the databases searched and the number of selected papers at each stage. Different approaches were followed to select grey and academic literature.
@@ 3.4.1. Academic Literature Selection
Searches in Scopus, IEEE Xplore, ACM DL, and Web of Science yielded 1,481 papers. After duplicate removal, 374 unique papers remained. Applying the inclusion and exclusion criteria, the selection was narrowed to 47 relevant papers. An additional search was conducted using Google Scholar, which identified two more articles, bringing the total to 49 papers.
Figure 2. The MLR study selection process. Steps are shown sequentially for both academic and grey literature. After final screening, inter-rater reliability was assessed using Cohen’s Kappa (Landis and Koch, 1977), yielding $\kappa=0.826$ for academic and $\kappa=0.857$ for grey literature, indicating strong reviewer agreement.
@@
Category Questions Criteria Score
@@
2/2
Date (Q6). Is the publication date of the source within the past eight years? 1/1
@@ 3.4.2. Grey Literature Selection
We followed the approach of other multi-vocal and grey literature reviews  (Garousi and Mäntylä, 2016; Garousi et al., 2017; Islam et al., 2019; Butijn et al., 2020; Garousi et al., 2019) and used Google Search to gather grey literature. As described by these reviews, we examined the first ten pages of results, as this was deemed sufficient to identify the most pertinent literature since Google’s algorithm emphasizes the most relevant results on the first pages  (Soldani et al., 2018; Verdecchia et al., 2019; Islam et al., 2019). We continued scanning further only if necessary, stopping when no relevant new articles appeared (Islam et al., 2019). To avoid bias, we searched in incognito mode, following the recommendations of Rahman et al. (Rahman et al., 2023). We assessed the quality and relevance of the primary grey literature sources we obtained, as it cannot be assumed that their quality is guaranteed. The quality assessment criteria suggested by Garousi et al(Garousi and Mäntylä, 2016) have been used for this purpose (see Table 5). Those criteria comprise three quality categories, ranging from the authority of the producer to the publication date (see the first column of Table 5). These categories encompass six criteria (see the second column) that were assessed individually for each grey literature item. The evaluation was conducted independently by the first two authors of this study, who indicated whether each item (a) met or (b) failed to meet a criterion. In cases where a criterion could not be assessed, such as when author or affiliation details were missing, it was treated as not satisfied. The final column of Table 5 specifies how many of the criteria needed to be satisfied for inclusion. For example, an item had to meet at least 2 out of 3 criteria related to the authority of the source to be considered valid. To further ensure internal agreement and consistency in applying these quality checks, all disagreements between reviewers were resolved through discussion. As a result of this procedure, 47 grey literature sources were selected. These, together with 49 academic papers, resulted in a total of 96 sources used for data extraction and synthesis (Figure 2). The distribution of selected sources across venues is summarized in Table 6.
@@
Websites & Blogs White Papers
(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Zhang et al., 2021a; Cotroneo et al., 2022; Camacho et al., 2022; Torkura et al., 2021; Poltronieri et al., 2022; Zhang et al., 2023; Nikolaidis et al., 2021; Frank et al., 2021b, a; Pierce et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Chen et al., 2024; Nikolaidis et al., 2023; Ma et al., 2023; Meiklejohn et al., 2021; Frank et al., 2023; Sondhi et al., 2021; Kassab et al., 2022; Jernberg et al., 2020; Konstantinou et al., 2021; Pulcinelli et al., 2023; Ji et al., 2023; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Park et al., 2019; Klinaku et al., 2022; Rivera et al., 2023; Camilli et al., 2022; Yu et al., 2021; Siwach et al., 2022; Basiri et al., 2016; Tucker et al., 2018; Alvaro and Tymon, 2017; Sousa et al., 2018; Alvaro et al., 2016) = 49 (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Hawkins, 2020; Nombela, 2023; Le, 2022; National Australia Bank, 2020; Shah, 2021; Krishnamurthy, 2021; Moon, 2022; Katirtzis, 2022; Wachtel, 2023; Durai, 2022; Anwar and Arunachalam, 2019; Tomka, 2024; Dua, 2024; Bocetta, 2019; Mooney, 2023; Wu, 2021; Lardo, 2019; Miles, 2019a; Bhaskar, 2022; Azure Readiness, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Sbiai et al., 2023; Abdul, 2024; Roa, 2022; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Vinisky, 2024; Colyer, 2019; Hochstein and Rosenthal, 2016; Butow, 2018; Wilms, 2018) = 44 (Palani and Gupta, 2023a; Tivey et al., 2019; Starr et al., 2022) = 3